AB INITIO CALCULATIONS OF ASTRONOMICALLY INTERESTING ANIONS

\underline{HARSHAL GUPTA}, Harvard-Smithsonian Center for Astrophysics, and Division of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138, and Institute for Theoretical Chemistry, Department of Chemistry and Biochemistry, The University of Texas at Austin, Austin, TX 78712, USA; JOHN F. STANTON, Institute for Theoretical Chemistry, Department of Chemistry and Biochemistry, The University of Texas at Austin, Austin, TX 78712; PATRICK THADDEUS, Harvard-Smithsonian Center for Astrophysics, and Division of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138, USA.

The recent laboratory detections of the linear carbon chain anions, C_{2n}H^{−} (n = 1, 2, 4), in the radio band were supported by high level ab initio calculations of their rotational constants. The methods employed in these calculations and the accuracy attained (∼ 0.1%) will be discussed. Our ongoing theoretical work towards guiding radio searches for other molecular anions will be described.